# Joule-Thomson expansion for quantum corrected AdS-Reissner-Nördstrom black holes in Kiselev spacetime with Barrow fractal entropy

### v1

@@ IV Numerical results
Figure 8: This picture shows the isenthalpic curves as a function of the pressure as solid lines with fixed values of $M=1.8$ (upper left panel), $M=2.0$ (upper right panel), $M=2.5$ (lower left panel), and $M=2.9$ (lower right panel), with fixed $\omega=-1$, $a=0.1$, $Q=1.0$, and $c=0.1$, for different values of the correction in entropy: $\Delta=0$ (blue line), $\Delta=1/2$ (red line), $\Delta=1$ (black line). Note that these curves have common zeros independent of the $\Delta$ values. Inversion temperature curves from Eq. (25) are represented by dashed lines for the same choice of parameters. These panels also show a transition behavior around $M=2.5$. For lower masses ($M<2.5$, upper panels) the increase of $\Delta$ decreases the isenthalpic curves, and for $M\geq 2.5$ the increase of $\Delta$ increase the isenthalpic curves.
@@
Figure 9: Isenthalpic curves, obtained with the parameters $\omega=-1$, $a=0.1$, $Q=1.0$, $c=0.1$, $M=2.0$. Here, we show that the two zeros in the $T\times P$ plane for the isenthalpics are identical for a given fixed mass $M$ and independent, as expected, of $\Delta$.
@@
Figure 10: Isenthalpic curves, obtained with the parameters $\omega=-1$, $a=0.1$, $Q=1.0$, $c=0.1$, $M=2.0$. Here, we show that the curves intersect at different values between the zeros.
@@
from which one sees that these zeros are independent of $\Delta$. Using the relation between the BH mass and the radius of the external horizon, Eq. (17), with a given fixed mass, we find an equation for $r_{+}$, the roots are $r_{+1}=0.100001$, $r_{+2}=0.352294$, $r_{+3}=5.64764$, for $M=2$, $\omega=-1$, $a=0.1$, $Q=1.0$, and $c=0.1$. The first root $r_{+1}$ implies a negative pressure $P_{1}=-716$, so it is not physically realizable. The other two roots imply the pressures $P_{2}=2.26$, and $P_{3}=0.0107$, which correspond to the values seen in Figs. 8 and 9 for $T=T_{i}=0$.
@@ V Conclusions and final considerations
In this work, we have investigated the effect of the modification of Barrow-Bekenstein-Hawking black hole entropy on the Joule-Thomson inversion temperature and isenthalpic curves. We saw that choosing different values of the Barrow fractal parameter $\Delta$ produces modifications in general, reducing the JT inversion temperature and increasing the corresponding pressure. This was done in the context of the Kiselev spacetime with quantum corrections.
After obtaining the general equations for the mass of BH, Eq. (17), the temperature of the Barrow-Bekenstein-Hawking BH, Eq. (21), and the Joule-Thomson inversion temperature, Eq. (25), we proceed to a numerical study of the problem, fixing some parameters and varying others. In particular, in Fig. 1, we considered the JT inversion temperature as a function of the pressure for fixed $\omega$, $c$, $a$, $Q$, varying the Barrow fractal parameter $\Delta$: 0, 1/4, 1/2, 3/4 and 1. In this picture, it is clearly seen that the increase in the values of $\Delta$ decreases the inversion temperature and increases the pressure. This picture also shows that the $T_{i}\times P$ curves cross for different values of $\Delta$.
In Fig. 2, we examined the behavior of the curves $T_{i}\times P$, fixing the parameters $\omega$, $c$, $a$, while varying the values of the BH charge $Q$: 1, 2, 3, and 6. Apart from a nontrivial characteristic of these curves at very low temperature and pressure (left panel), the general profile (right panel) is that increasing the values of $\Delta$ reduces $T_{i}$ increasing the pressure. The effect of increasing the BH charge is the opposite: it increases $T_{i}$ and decreases $P$. In addition, the increase of $\Delta$ tends to yield more straight curves than that of low values of $\Delta$. This is basically the same behavior found in Figs. 3-7, where we vary different parameters and the values of $\Delta$.
@@
It is also interesting to mention that for the various Kiselev scenarios described by the examined values of $\omega$, the effect of the Barrow parameter $\Delta$ is essentially the same, as can be seen in Figs. 3 and 5-7. Despite some particular behavior for the small temperature and pressure regime, for large values of these quantities, as the values of $\Delta$ increase, the slope of the $T_{i}$ curve decreases, implying an increase in pressure. This can be understood since $T=(\partial M/\partial S)_{P,Q}$, Eq. (19), so that increasing the entropy $S$, should decrease $T$.
Of special mention is the case of Fig. 4 where we saw clearly the difference between both manifestations of quantum corrections, the one associated with parameter $a$ inside the metric, Eq. (15), and the other related to $\Delta$ appearing in the expression of entropy of the so-called Barrow entropic model, Eq. (18). The metric quantum correction deals with the geometry of the spacetime fabric and, on the other hand, the entropic quantum correction is related to the storage of the information in the event horizon which reflects a non smooth surface, a fractal one. The latter is connected to the quantum microstates of the BH. Given the different nature of these corrections, we expect a difference in the inversion temperature curves when we vary those parameters. The variation of $\Delta$ changes the inclination of the curves and the variation of the $a$-parameter dislocated the initial points of each curve on the pressure scale; this is well defined in Fig. 4.

### v2

@@ IV Numerical results
Figura 8: This picture shows the isenthalpic curves as a function of the pressure as solid lines with fixed values of $M=1.8$ (upper left panel), $M=2.0$ (upper right panel), $M=2.5$ (lower left panel), and $M=2.9$ (lower right panel), with fixed $\omega=-1$, $a=0.1$, $Q=1.0$, and $c=0.1$, for different values of the correction in entropy: $\Delta=0$ (blue line), $\Delta=1/2$ (red line), $\Delta=1$ (black line). Note that these curves have common zeros independent of the $\Delta$ values. Inversion temperature curves from Eq. (25) are represented by dashed lines for the same choice of parameters. These panels also show a transition behavior around $M=2.5$. For lower masses ($M<2.5$, upper panels) the increase of $\Delta$ decreases the isenthalpic curves, and for $M\geq 2.5$ the increase of $\Delta$ increase the isenthalpic curves.
@@
Figura 9: Isenthalpic curves, obtained with the parameters $\omega=-1$, $a=0.1$, $Q=1.0$, $c=0.1$, $M=2.0$. Here, we show that the two zeros in the $T\times P$ plane for the isenthalpics are identical for a given fixed mass $M$ and independent, as expected, of $\Delta$.
@@
Figura 10: Isenthalpic curves, obtained with the parameters $\omega=-1$, $a=0.1$, $Q=1.0$, $c=0.1$, $M=2.0$. Here, we show that the curves intersect at different values between the zeros.
@@
from which one sees that these zeros are independent of $\Delta$. Using the relation between the BH mass and the radius of the external horizon, Eq. (17), with a given fixed mass, we find an equation for $r_{+}$, the roots are $r_{+1}=0.100001$, $r_{+2}=0.352294$, $r_{+3}=5.64764$, for $M=2$, $\omega=-1$, $a=0.1$, $Q=1.0$, and $c=0.1$. The first root $r_{+1}$ implies a negative pressure $P_{1}=-716$, so it is not physically realizable. The other two roots imply the pressures $P_{2}=2.26$, and $P_{3}=0.0107$, which correspond to the values seen in Figs. 8 and 9 for $T=T_{\rm i}^{\rm JT}=0$.
@@ V Conclusions and final considerations
In this work, we have investigated the effect of the modification of Barrow on Bekenstein-Hawking black hole entropy, speccifically on the Joule-Thomson inversion temperature and isenthalpic curves in AdS-Reissner-Nordström black holes in Kiselev spacetime with quantum corrections. We saw that choosing different values of the Barrow fractal parameter $\Delta$ produces modifications in general, reducing the JT inversion temperature and increasing the corresponding pressure.
After obtaining the general equations for the mass of the BH, Eq. (17), its temperature, Eq. (21), and the Joule-Thomson inversion temperature, Eq. (25), we proceed to a numerical study of the problem, fixing some parameters and varying others. In particular, in Fig. 1, we considered the JT inversion temperature as a function of the pressure for fixed $\omega$, $c$, $a$, $Q$, varying the Barrow fractal parameter $\Delta$: 0, 1/4, 1/2, 3/4 and 1. In this picture, it is clearly seen that the increase in the values of $\Delta$ decreases the inversion temperature and increases the pressure. This picture also shows that the $T_{\rm i}^{\rm JT}\times P$ curves cross for different values of $\Delta$.
In Fig. 2, we examined the behavior of the curves $T_{\rm i}^{\rm JT}\times P$, fixing the parameters $\omega$, $c$, $a$, while varying the values of the BH charge $Q$: 1, 2, 3, and 6. Apart from a nontrivial characteristic of these curves at very low temperature and pressure (left panel), the general profile (right panel) is that increasing the values of $\Delta$ reduces $T_{\rm i}^{\rm JT}$ increasing the pressure. The effect of increasing the BH charge is the opposite: it increases $T_{\rm i}^{\rm JT}$ and decreases $P$. In addition, the increase of $\Delta$ tends to yield more straight curves than that of low values of $\Delta$. This is basically the same behavior found in Figs. 3-7, where we vary different parameters and the values of $\Delta$.
@@
It is also interesting to mention that for the various Kiselev scenarios described by the examined values of $\omega$, the effect of the Barrow parameter $\Delta$ is essentially the same, as can be seen in Figs. 3 and 5-7. Despite some particular behavior for the small temperature and pressure regime, for large values of these quantities, as the values of $\Delta$ increase, the slope of the $T_{\rm i}^{\rm JT}$ curve decreases, implying an increase in pressure. This can be understood since $T=(\partial M/\partial S)_{P,Q}$, Eq. (19), so that increasing the entropy $S$, should decrease $T$.
There are various types of quantum corrections in gravity which depend on the specific situation or energy and length scales. In particular, in the case of Fig. 4, we clearly saw the numerical difference between the manifestations of the quantum corrections studied here, the one associated with parameter $a$ inside the metric, Eq. (15), and the other related to $\Delta$ appearing in the expression of entropy of the so-called Barrow entropic model, Eq. (18). The metric quantum correction deals with the geometry of the spacetime fabric and, on the other hand, the entropic quantum correction is related to the storage of the information in the event horizon which reflects a non smooth surface, a fractal one. The latter is connected to the quantum microstates of the BH. The variation of $\Delta$ changes the inclination of the curves and the variation of the $a$-parameter dislocated the initial points of each curve on the pressure scale. Since these corrections appear clearly in Fig. 4, one can see that both perturbations are consistent and with similar energy or length scales for the chosen values of the parameters.